Davis Inc's bonds currently sell for $800 and have a par value of 1000. They pay a $50 annual coupon and have a 20 year maturity, but they can be called in 5 years at $1200. What is their Capital Gain Yield (CGY)?
Answer #1

Solution :-

Here the question specifically ask about Capital Gain Yield not the total yield means not including current yield

Capital Gain Yield = Means not include the Annual Coupon income

Capital Gain Yield = ( P1 - P0 ) / P0 = ( 1200 - 800 ) / 800 = 400 / 800 = 50%

Which means 50% capital gain yield in 5 years

If we want to know annual capital gain yield , then it equals to ( 1 + 50% )1/5 - 1  = ( 1.50 )1/5 - 1 = 1.08447 = 8.447%

If there is any doubt please ask in comments